I need help with this question. Find the balance for $1,000 invested at a rate of 8% for four years if the interest is compunded quarterly.
Answer by lwsshak3(11628)

I need help with this question. Find the balance for $1,000 invested at a rate of 8% for four years if the interest is compunded quarterly
..
Compound Interest Formula:
A=P(1+r)^n
P=Initial investment
r=interest rate/period
n=number of periods
A=Investment worth after n periods
..
For given problem,
P=1000
r=.08/4=.02
n=8*4=32
A=1000(1+.02)^32
A=$1884.54
ans:
$1000 invested at an annual interest rate of 8% compounded quarterly would be worth $1884.54 after 8 years